# Neural Network Enumerator for Atmospheric Chemical ODE

### v1

@@ 3.2.1 Time embedding and Implicit Neural Representation (INR) mapping
To obtain implicit neural representation (INR), we apply sine as a periodic activation function to the chemical input and hidden features, $x_{i}\rightarrow\phi_{i}(x_{i})=sin(W_{i}x_{i}+b),\text{where\ }W_{i}\sim\mathcal{U}(-6/\sqrt{(}N),6/\sqrt{(}N)),c\in\mathbb{R}$, as they are continuous changes in the physical world and it is memory efficient to compute the derivatives in the sine space without being constrained by discrete data samples. The initialized weight parameters ensure that the input to each sine activation is normally distributed with a standard deviation of 1. Physically, we can interpret the weights of the INR function as angular frequencies while the biases are phase offsets. Applying the sine function keeps signal amplitude constant while expanding frequency bands for high-frequency modeling.
@@ 3.2.2 Attention for chemical representation
Attention has been widely used in image, video, and language processing. Its success comes from its efficient nonlocal feature representation. For chemical modeling, we also expect an efficient approach that can model the long-term time evolution, so that similar chemical behavior across time or different chemical compounds can be utilized for pattern matching. In order to achieve that, we propose a time-dependent attention module, which can encode and decode the chemical data to the latent space for implicit neural representation. Mathematically, we can define the process as,
$z_{\lambda}(t)=z_{\lambda}(t)+\text{ffn}\left(z_{\lambda}(t)+\sigma\left(\frac{Q_{z}K_{z}^{T}}{\sqrt{(}d)}\right)V_{z}\right)$ (3)
where $z_{\lambda}(t)=sin\left[W(z_{0})+b\right]+\lambda(t)$, $Q_{z}=z_{\lambda}(t)W_{Q}$, $K_{z}=z_{\lambda}(t)W_{K}$, and $V_{z}=z_{\lambda}(t)W_{V}$. The $W_{*}$ is the learnable parameter. $z_{0}$ is the output feature of the INR mapping layer. As depicted in Figure 1, given the latent space chemical features, we apply the INR mapping first. Then we use the attention module to learn the nonlocal correlations. A feedforward network (ffn) is used to learn the residues for network update. Note that the activation of ffn is also a sine function to preserve the periodic behavior of chemical reactions.
@@ 3.2.3 Fourier Neural Operator (FNO)
To explore the derivatives of the chemical reactions, introducing ODE in the neural network can preserve the underlying physical laws. In recent works Chen et al. (2018); Li et al. (2023, 2020), neuralODE, like Torchdiffeq (Chen et al. (2018)), has been used for GPU based ODE approximation. However, the disadvantage is that it is not numerically stable and rather slow in computation. We propose to further simplify the ODE process as a Neural Fourier Operation. Given the continuous time representation, we use Fourier Transformation to project the signal to the frequency domain and learn the parametric partial differentiation process.
@@
The advantages of using FNO for ODE computation are 1) it is faster to use Fourier Transform than convolution, as it is quasilinear, where the full standard integration of n points has complexity $O(n^{2})$; 2) the input and outputs of PDEs are continuous functions, so it is efficient to represent them in the Frequency domain for global convolution. In the experiments, we would show the complexity comparisons to demonstrate the efficiency of FNO.
3.2.4 Physical-informed losses.
@@ 3.2.5 Evaluation
Furthermore, we also calculate the running time of the proposed model to compare with numerical simulation to see its efficiency, which could indicate how it can be utilized for large-scale modeling.
@@ 4.1 Implementing Details
    Task 1: Intraperiod chemical prediction. Given the same 49 chemical compounds, we take their initial values ($x_{0}^{i},i=1,2,...,49$) and environmental parameters $k$ to predict their future states. We use the ARCA box to simulate one-hour changes in chemical concentration and sample the observations every 5 minutes. The objective of the ChemNNE is to predict the results of all 11 time steps, as $x^{i}_{t},i=1,2,...,49,t=1,2,...,11$.
@@
    Task 2: Interperiod chemical prediction. Based on the simulated concentrations of ARCA (concentrations above a defined threshold), we select 100 significant chemical compounds that have high impacts on air quality. We train the proposed model to take the same 49 chemical values ($x^{i}_{0},i=1,2,...,49$) and environmental parameters $k$, and predict the time evolutions of the selected 100 chemical compounds ($y^{i}_{t},i=1,2,...,100,t=1,2,...,11$).
@@
    Task 3: Hybridperiod chemical prediction. Similarly, as Task 2, we further select 300 more new chemical compounds that are generated over time. This is a much more challenging task. We design the model to take 49 initial chemical values and 100 new chemical values from Task 2, to output another 300 new chemical compounds at different time steps as ($z^{i}_{t},i=1,2,...,400,t=1,2,...,11$).
@@
To standardize the data for neural network training, we first take the logarithm of true chemical output to the base 10, then we normalize all data by dividing the maximum value, approximately 31.5. For the environmental parameters, we take the maximum and minimum values to normalize them to [-1, 1]. To further increase the data variety, we apply data augmentation to the training set. Specifically, we randomly roll the time evolution of the observation as $\hat{y}_{i}^{t}=y_{i}^{t+\tau},i=1,2,...,100,t=1,2,...,11$, where
$\hat{y}_{i}^{t}=\left\{\begin{array}[]{rcl}y_{i}^{t+\tau}&\mbox{for}&\tau<11-t\\ y_{i}^{t+\tau-11}&\mbox{for}&\tau>11-t\end{array}\right.\text{where }i=1,2,...,t=1,2,...,11$ (7)
$\bullet$ Parameter setting. We train ChemNNE using Adam optimizer with the learning rate of $1\times 10^{-3}$. The batch size is set to 4096 and ChemNNE is trained for 100k iterations (about 2 hours) on a PC with one NVIDIA V100 GPU using PyTorch deep learning platform. The weighting factors in the total loss are defined empirically as: $\alpha_{1}=1,\alpha_{2}=10,\alpha_{3}=10,\alpha_{4}=1,\alpha_{5}=0.001$.
Figure 3: Visualization of the time evolution of chemistry in Task 1. In (a), (b), and (c), we show the ground truth as dashed black lines and the predictions as solid lines with different colors. We also enlarge the region in red boxes to highlight the prediction errors.
@@
Figure 4: Visualization of the time evolution of chemistry in Task 2 and 3. In (a), and (b), we show the ground truth as dashed black lines and the predictions as solid lines with different colors. Task 2 predicts 100 unknown chemical components and Task 3 predicts 400 unknown chemical components.
4.2 Overall comparison with state of the arts
To demonstrate the efficiency of our proposed ChemNNE , we compare it with two state-of-the-art methods: UNet (Williams et al. (2024)), NeuralODE (Chen et al. (2018)), which are widely used in neural ODE/PDE processes. The comparison is shown in Table 2. We compare different approaches on both validation and testing datasets in logarithm scales. We also show the model complexity by running time (seconds) and number of Multiply–accumulate operations (MACs). Compared to NeuralODE and UNet, we can observe that using our proposed ChemNNE can achieve the best performance in terms of RMSE, MAE and MBE. For instance, ours can improve the RMSE by about 0.28 $\sim$ 0.41 in Task 1, 0.1$\sim$ 0.3 in Task 2, and 0.14$\sim$ 0.23 in Task 3. Comparing different tasks, we can see that the improvements from Task 1 to Task 3 are reduced which also indicates that the modeling difficulty is increased when more unknown chemical components are estimated. From the computation complexity, ours has a comparable number of operations as the other two and achieves a similar running speed. As the baseline, the numerical chemical model takes 0.05 seconds to run one simulation. It indicates that ours can be used as a fast enumerator to accelerate the chemical modeling process. For visualization, we take different results on Task 1 and visualize them in Figure 3. We enlarge the red-box regions to demonstrate better the differences between ground truth (dashed black lines) and predictions (solid color lines). We can see that using ours can accurately align with the ground truth. For Task 2 and 3, we show the results in Figure 4. We can see that the color lines are overlapped with the dashed black lines, which indicates that our predictions align well with the ground truth values.
@@ 4.2 Overall comparison with state of the arts
AE Attn Time emb INR FNO RMSE MAE MBE RMSE MAE MBE RMSE MAE MBE
@@ 4.3 Ablation studies
Figure 5: Training loss comparison among ours and others. We show different approaches in different colors, and we can see the improvements of using our proposed ChemNNE in both convergence speed and final loss.
We use the validation dataset of all three Tasks to demonstrate the effects of different key components, including AE (AutoEncoder), attention (attn), time embedding (Time emb), INR, FNO, and their combinations. Using AE is our baseline. We can see that the last row is our final model, which shows the best performances in all metrics. Individually, we can see that using attention and FNO can achieve the most improvements by approximately 0.1 $\sim$ 0.4 in terms of RMSE. Using time embedding and INR can also improve the RMSE by about 0.01 $\sim$ 0.2.
To further demonstrate the effect of the key components of attention and FNO, we visualize the training loss convergences in Figure 5. Compared to Unet and NeuralODE, we can see that ours with attention and/or FNO achieves the fastest convergence speed and lowest training loss.
@@
Finally, we train our proposed ChemNNE with different loss combinations, so that we can validate the effects of using physical-informed losses for model optimization. As the baseline, we train the model with MSE loss ($L_{recon}$), and then we individually add other losses to retrain the model, including first- and second-order derivative losses (Derivs), identity loss (IDN), and mass conservation loss (Mass). In Table 4, we show the RMSE, MAE, and MBE results on three tasks. Our observations are: 1) from columns 2 and 3, we can see that using identity and mass conservation losses have a more significant loss drop, about 0.03 $\sim$ 0.1 in RMSE, 2) using derivative loss has more visible effects on Task1 than Task 2 and 3 because Task 1 is intraperiod chemical prediction, that is, the derivative can better constrain the chemical reaction process for accurate estimation.
@@ 4.4 Statistical analysis
To study in depth the proposed ChemNNE on the ability of chemical modeling, we study the error distributions among different chemical components and time steps, so that we can see whether the model can estimate the patterns for each individual component, as well as long-term regression. In Figure 6, we show the error distribution in Task 1. We can see that 1) the model performs unevenly on different chemical components in terms of means and variances, and 2) the errors increase when predicting further steps in the future.
For Tasks 2 and 3, we have a much larger number of chemical components, 100 to 400 chemical components, for computation. To efficiently demonstrate the model performance, we show the 2D heatmap of mean and variance distributions on the validation datasets. Figure 7 and Figure 8 show the mean and variance of error distributions in all three tasks. The brighter the color, the higher the errors the model produces. From Figure 7, we can observe that the average performance differences depend on the specific chemical components. From Figure 8, we can see that the error variances vary because of the time steps. The farther the future step to predict, the higher the variances the model gets. The patterns also match our observations in Figure 6.
@@ 5 Conclusion
In this paper, we propose the first work on a neural network enumerator for fast chemical modeling, dubbed ChemNNE . We utilize the numerical simulation to generate a large-scale chemical dataset which is used to train the proposed ChemNNE , so that it can learn the hidden inter- and intra-correlations among chemical molecules. We propose to combine attention and neural ODE operations to model the time-dependent chemical reactions. Meanwhile, we lift the neural ODE as a Fourier domain convolution such that we can efficiently model the global continuous ODE. The time embedding and INR operators also help model the oscillation patterns to mimic the concentration changes of chemistry. To demonstrate the efficiency and effectiveness of the proposed model, we test it on three tasks, including intraperiod, interperiod and hybridperiod chemical predictions. Extensive experiments show that ours achieves the best performance in both accuracy and running speed. This work paves a new direction in AI for atmospheric chemistry, and we will continue to explore graph neural networks, equivalent neural networks, and other advanced models to integrate chemical knowledge for physics-informed processing.

### v3

@@ 3.2.2 Attention for chemical representation
Attention has been widely used in image, video, and language processing. Its success comes from its efficient nonlocal feature representation. For chemical modelling, we also expect an efficient approach that can model the long-term time evolution, so that similar chemical behavior across time or different chemical compounds can be utilized for pattern matching. In order to achieve that, we propose a time-dependent attention module, which can encode and decode the chemical data to the latent space for implicit neural representation. Mathematically, we can define the process as,
$z_{\lambda}(t)=z_{\lambda}(t)+\text{ffn}\left(z_{\lambda}(t)+\sigma\left(\frac{Q_{z}K_{z}^{T}}{\sqrt{d}}\right)V_{z}\right)$ (3)
where $z_{\lambda}(t)=sin\left[W(z_{0})+b\right]+\lambda(t)$, $Q_{z}=z_{\lambda}(t)W_{Q}$, $K_{z}=z_{\lambda}(t)W_{K}$, and $V_{z}=z_{\lambda}(t)W_{V}$. $\sigma$ is the softmax function. The $W_{*}$ is the learnable parameter. $z_{0}$ is the output feature of the Sinusoid mapping layer. $d$ is the dimension of the learned feature vector. As depicted in Figure 1, given the latent space chemical features, we apply the Sinusoid mapping first. Then we use the attention module to learn the nonlocal correlations. A feedforward network (ffn) is used to learn the residues for network update. Note that the activation of ffn is also a sine function to preserve the periodic behavior of chemical reactions.
@@ 3.2.3 Fourier Neural Operator (FNO)
To explore the derivatives of the chemical reactions, introducing ODE in the neural network can preserve the underlying physical laws. In recent works Chen et al. (2018); Li et al. (2023, 2020), neuralODE, like Torchdiffeq (Chen et al. (2018)), has been used for GPU based ODE approximation. However, the disadvantage is that it is not numerically stable and rather slow in computation. We propose to further simplify the ODE process as a Neural Fourier Operation. Given the discrete-time representation across multiple chemical components, we use Fourier Transformation to project the signal to the frequency domain and learn the multi-grid nonlinear interactions across temporal and component domains.
@@
The advantages of using FNO for ODE computation are 1) CNN and transformers can only learn on fixed-scale grids and fail to capture the fine scales of multi-scale systems, while FNO is grid-invariant since the operations outside of the Fourier layers act point-wise on the spatial domain, 2) it is faster to use Fourier Transform than convolution, as it is quasilinear, where the full standard integration of n points has complexity $O(n^{2})$; 3) the input and outputs of PDEs are discrete functions, so it is efficient to represent them in the Frequency domain for global convolution. In the experiments, we would show the complexity comparisons to demonstrate the efficiency of FNO.
3.2.4 Physics-informed losses.
@@ 3.2.5 Evaluation
Furthermore, we also calculate the running time of the proposed model to compare with numerical simulation to see its efficiency, which could indicate how it can be utilized for large-scale modelling.
@@ 4.1 Implementation Details
    Task 1: Intraperiod chemical prediction. Given the same 49 chemical compounds, we take their initial values ($x_{0}^{i},i=1,2,...,49$) and environmental parameters $k$ to predict their future states. We use the ARCA box to simulate one-hour changes in chemical concentration and sample the observations every 5 minutes. The objective of the ChemNNE is to predict the results of all 12 time steps, as $x^{i}_{t},i=1,2,...,49,t=1,2,...,12$.
@@
    Task 2: Interperiod chemical prediction. Based on the simulated concentrations of ARCA (concentrations above a defined threshold), we select 100 significant chemical compounds that have high impacts on air quality. We train the proposed model to take the same 49 chemical values ($x^{i}_{0},i=1,2,...,49$) and environmental parameters $k$, and predict the time evolutions of the selected 100 chemical compounds ($y^{i}_{t},i=1,2,...,100,t=1,2,...,12$).
@@
    Task 3: Hybridperiod chemical prediction. Similarly, as Task 2, we further select 300 more new chemical compounds that are generated over time. This is a much more challenging task, and we need additional chemical information to help the model understand the chains of chemical reactions. We design the model to take 49 initial chemical values and 100 new chemical values from Task 2, to output another 300 new chemical compounds at different time steps as ($z^{i}_{t},i=1,2,...,300,t=1,2,...,12$).
@@
To standardize the data for neural network training, we first take the logarithm of true chemical output to the base 10, then we normalize all data by dividing the maximum value, approximately 31.5. For the environmental parameters, we take the maximum and minimum values to normalize them to [-1, 1]. To further increase the data variety, we apply data augmentation to the training set. Specifically, we randomly roll the time evolution of the observation as $\hat{y}_{i}^{t}=y_{i}^{t+\tau},i=1,2,...,100,t=1,2,...,12$, where
$\hat{y}_{i}^{t}=\left\{\begin{array}[]{rcl}y_{i}^{t+\tau}{}{}{}{}{}&\mbox{for}&\tau<12-t\\ y_{i}^{t+\tau-12}&\mbox{for}&\tau>12-t\end{array}\right.\text{where }i=1,2,...,t=1,2,...,12$ (7)
$\bullet$ Parameter setting. We train ChemNNE using Adam optimizer with the learning rate of $1\times 10^{-3}$ and halved around 20K iterations. The batch size is set to 4096 and ChemNNE is trained for 100k iterations (about 2 hours) on a PC with one NVIDIA V100 GPU using PyTorch deep learning platform. The weighting factors in the total loss are defined empirically as: $\alpha_{1}=1,\alpha_{2}=10,\alpha_{3}=10,\alpha_{4}=1,\alpha_{5}=0.001$.
Figure 3: Visualization of the time evolution of chemistry in Task 1, 2 and 3. In (a), (b), and (c), we show the ground truth as red lines and the predictions with different colors. We pick four different chemical compounds for comparison, and we also enlarge the region in red boxes to highlight the prediction errors.
@@
Figure 4: Visualization of the mean errors of the time evolution in Task 1, 2 and 3. We average all chemical compounds and show the mean absolute errors between ground truth and other predictions across different time steps.
4.2 Overall comparison with state-of-the-arts
To demonstrate the efficiency of our proposed ChemNNE , we compare it with two state-of-the-art methods: UNet (Williams et al. (2024)), NeuralODE (Chen et al. (2018)), which are widely used in neural ODE/PDE processes. The comparison is shown in Table 2. We compare different approaches on both validation and testing datasets in logarithm scales. We also show the model complexity by running time (seconds) and number of Multiply–accumulate operations (MACs). Compared to NeuralODE and UNet, we can observe that using our proposed ChemNNE can achieve the best performance in terms of RMSE, MAE and MBE. For instance, ours can improve the RMSE by about 0.28 $\sim$ 0.41 in Task 1, 0.1$\sim$ 0.3 in Task 2, and 0.14$\sim$ 0.23 in Task 3. Comparing different tasks, we can see that the improvements from Task 1 to Task 3 are reduced which also indicates that the modelling difficulty is increased when more unknown chemical components are estimated. From the computation complexity, ours has a comparable number of operations as the other two and achieves a similar running speed. As the baseline, the numerical chemical model takes 0.05 seconds to run one simulation. It indicates that ours can be used as a fast emulator to accelerate the chemical modelling process. In practical applications, neural surrogate models can detect and respond to subtle climate changes far more quickly than traditional numerical models. This rapid detection has the potential to significantly reduce human and economic costs associated with catastrophic climate events. For visualization, we take different results on Tasks 1, 2, and 3 and visualize them in Figure 3. In tasks 1, 2, and 3, there are over 100 different chemical compounds, which makes it difficult to visualize them all in one figure. We randomly pick 4 individual chemical compounds and draw them in different line styles (solid, dotted, dashed, dash-dot). For each compound, we show the predictions of UNet, NeuralODE and ours in green, blue and red color. The ground truth value is shown in black color. We enlarge the red-box regions to demonstrate better the differences between ground truth and predictions. We can see that using ours can accurately align with the ground truth, while UNet and NeuralODE still produce distinct gaps between the ground truth. For a complete chemical compound comparison, you can find the figures in the supplementary material.
Meanwhile, we calculate the mean absolute prediction errors across different chemical compounds and summarize errors by time steps. In Figure 4, we compare different approaches to three different tasks. We can see that ours achieves the lowest errors across different time steps in all tasks. We can also observe that the further time steps, the higher the errors we get, which fits our assumption that long-term chemical prediction produces higher uncertainty. Depending on the difficulty of different tasks, we can also see that task 2 has the overall highest errors, and task 1 has the lowest errors. It is interesting to note that Task 3 uses 100 chemical compounds actively involved in the chemical reactions as additional input. This allows the model to better understand the correlations between the initial chemical values and the 300 new chemical compounds. For instance, knowing the concentration of OH (one of the 100 significant chemical compounds for air quality) can help predict the concentration changes of C7300H and OXYLOOH (two of the 300 predicted chemical compounds). As a result, the model for Task 3 performs slightly better than that for Task 2.
@@ 4.2 Overall comparison with state-of-the-arts
AE Attn Time emb Sinusoid FNO RMSE MAE MBE RMSE MAE MBE RMSE MAE MBE
@@ 4.3 Ablation studies
Figure 5: Training loss comparison among ours and others. We show different approaches in different colors. We enlarge the red-boxed region and display it at the upper center. We can see the improvements of using our proposed ChemNNE in both convergence speed and final loss.
We use the validation dataset of all three Tasks to demonstrate the effects of different key components, including AE (AutoEncoder), attention (attn), time embedding (Time emb), Sinusoid, FNO, and their combinations. Using AE is our baseline. We can see that the last row is our final model, which shows the best performances in all metrics. Individually, we can see that using attention and FNO can achieve the most improvements by approximately 0.1 $\sim$ 0.4 in terms of RMSE. Using time embedding and Sinusoid can also improve the RMSE by about 0.01 $\sim$ 0.2.
To further demonstrate the effect of the key components of attention and FNO, we visualize the training loss convergences in Figure 5. We enlarge the region within the red box and highlight it at the upper center to illustrate the loss differences. Compared to Unet and NeuralODE, we can see that ours with attention and/or FNO achieves the fastest convergence speed and lowest training loss.
@@
Finally, we train our proposed ChemNNE with different loss combinations, so that we can validate the effects of using physics-informed losses for model optimization. As the baseline, we train the model with MSE loss ($L_{recon}$), and then we individually add other losses to retrain the model, including first- and second-order derivative losses (Derivs), identity loss (IDN), and mass conservation loss (Mass). In Table 4, we show the RMSE, MAE, and MBE results on three tasks. Our observations are: 1) from columns 2 and 3, we can see that using identity and mass conservation losses have a more significant loss drop, about 0.03 $\sim$ 0.1 in RMSE, 2) using derivative loss has more visible effects on Task 1 than Task 2 and 3 because Task 1 is intraperiod chemical prediction, that is, the derivative can better constrain the chemical reaction process for accurate estimation.
@@ 4.4 Statistical analysis
To study in depth the proposed ChemNNE on the ability of chemical modelling, we study the error distributions among different chemical components and time steps, so that we can see whether the model can estimate the patterns for each individual component, as well as long-term regression. In Figure 6, we show the error distribution in Task 1. We can see that 1) the model performs unevenly on different chemical components in terms of means and variances, and 2) the errors increase when predicting further steps in the future.
For Tasks 2 and 3, we have a much larger number of chemical components, 100 to 300 chemical components, for computation. To efficiently demonstrate the model performance, we show the 2D heatmap of mean and variance distributions on the validation datasets. Figure 7 and Figure 8 show the mean and variance of error distributions in all three tasks. The brighter the color, the higher the errors the model produces. From Figure 7, we can observe that the average performance differences depend on the specific chemical components. From Figure 8, we can see that the error variances vary because of the time steps. The farther the future step to predict, the higher the variances the model gets. The patterns also match our observations in Figure 6.
@@ 5 Conclusion
In this paper, we propose the first work on a neural network emulator for fast chemical modelling, dubbed ChemNNE . We utilize the numerical simulation to generate a large-scale chemical dataset which is used to train the proposed ChemNNE , so that it can learn the hidden inter- and intra-correlations among chemical molecules. We propose to combine attention and neural ODE operations to model the time-dependent chemical reactions. Meanwhile, we lift the neural ODE as a Fourier domain convolution such that we can efficiently model the global continuous ODE. The time embedding and Sinusoid operators also help model the oscillation patterns to mimic the concentration changes of chemistry. To demonstrate the efficiency and effectiveness of the proposed model, we test it on three tasks, including intraperiod, interperiod and hybridperiod chemical predictions. Extensive experiments show that ours achieves the best performance in both accuracy and running speed. This work paves a new direction in AI for atmospheric chemistry, and we will continue to explore graph neural networks, equivalent neural networks, and other advanced models to integrate chemical knowledge for physics-informed processing.